# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 4.1 Strategy to reconcile the aging law and the slip law
However, a well-known side effect exists in its prediction of steady-state $\Phi_{\rm SS}(V)$, as pointed out by Kato & Tullis (2001) themselves (Fig. 6). As shown in Fig. 6b, $\Phi_{\rm SS}(V)$ for $V\ll V_{\rm c}$ is predicted to be higher than the experimental requirement (eq. 26) by $W(1)B$, where $W(1)=0.567...$ is the omega constant, contradicting the SS canon. This level shift of $\Phi_{\rm SS}$ around $V_{\rm c}$ results from the extra velocity-strengthening mechanism (i.e., the aging law) operating at $V\ll V_{\rm c}$, in addition to the slip law, which alone matches eq. (26) exactly. By fitting $\tau_{\rm SS}(V)$ data of Blanpied et al. (1998b),Kato & Tullis (2001) suggested the possible level shift at $V_{\rm c}=0.01\mu$m/s. However, the data have large scatterings and do not make strong evidence for the reality of the level shift. Furthermore, Kato & Tullis (2001) noted that $V_{\rm c}=0.01\mu$m/s contradicts behaviors in VS tests, as explained in the next paragraph.
Figure 6: Comparison of steady-state values predicted by the composite law (blue) and those given by the SS constraint (eq. 71; red). Plotted against the slip rate $V$ normalized by $V_{\rm c}$. (a) Steady-state values of $\Omega=V\theta/L$. (b) Steady-state values of strength $\Phi$. Difference of $\Phi_{\rm SS}(V)$ from the canonical value $\Phi_{*}-B\ln(V/V_{*})$, given by eqs. (11) and (71) with the reference slip rate $V_{*}$ taken to satisfy $V_{*}\gg V_{\rm c}$, is plotted after being normalized by $B$. $\Delta\Phi_{\rm SS}=\Phi_{\rm SS}(V)-\Phi_{\rm SS}(V_{\rm prior})$ denotes the variation in $\Phi_{\rm SS}$.
@@
Lastly, we briefly describe our strategy to reproduce the three canons. Among the three defects of the composite law mentioned above, the level shift of $\Phi_{\rm SS}(V)$, which violates the SS canon (eq. 26), is not directly rooted in the antinomy between the VS constraint (eq. 72) and SHS||||NSC constraint (eq. 73). Hence, the present development aims to predict $\Phi_{\rm SS}(V)$ exactly as eq. (26). In §4.2, we will show that it is possible to construct evolution laws that perfectly match the SS requirement (eq. 26) for all the ranges of $V$. Moreover, we will find functional forms to do so are fairly restricted. This finding has facilitated our development. Within such functional forms, we will try to narrow the conditions where the other two unavoidable deviations from canons during the strengthening phase of VS and SHS||||NSC tests arise. Eventually, we could see that our proposed evolution laws violate canons only at extreme conditions outside the conditions experimentally verified so far. As the constraints eqs. (72) and (73) are the slip law and the aging law for $\Omega\ll 1$, respectively, the present development, in which we purposefully control the unavoidable side effects of the aging-slip switch, squarely answers the long-standing problem of how to reconcile these two laws (Ampuero & Rubin, 2008).
Figure 7: Strength evolution in ideal VS tests predicted by the composite law. $L=D_{\rm c}$ and $\theta\gg\theta_{\rm X}$ are assumed. $V_{\rm before}$ is fixed at $100V_{\rm c}$. The visualizing scheme is the same as in Fig. 2. Solid lines of different colors show the strength evolutions for different values of $V_{\rm after}$. As a reference, the canonical behavior (eq. 32) for each $V_{\rm after}$ is shown by the dotted line of the same color, although not visible in cases with $V_{\rm after}\geq 0.1V_{\rm before}$ because it almost coincides with the prediction by the composite law.
@@ 4.2.1 Heuristic derivation
As $\psi$ is a rewritten form of $\phi$, $\psi$ has a unique zero at $\Omega=C_{\rm SS}$ [$\psi(\Omega)=0\Leftrightarrow\Omega=C_{\rm SS}$]. Figure 8 shows the behavior of evolution laws given by eq. (77) (or equivalently eq. 78). In the $V$-$\theta$ space, $\dot{\theta}$ varies only along the direction of the $V\theta$-axis while remaining constant along the direction normal to it (dotted lines in Fig. 8a).
@@
For $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, the expressions of $f_{t|{\rm NSC}}$ and $f_{\delta}V$ (eqs. 72 and 73) both fit the functional form eq. (78) we consider. As far as eq. (78) holds, the ratio $f_{t}/f_{t|{\rm NSC}}$ is also a function of $\Omega$, which we denote by $\psi_{\rm c}(\Omega)$ hereafter. Then, under eqs. (72), (73), and (78), the thermodynamically formulated $f$ (eq. 25) can be expressed as
@@
So far, eq. (77) is the only heuristic. Otherwise, the derivation down to eq. (83) from eqs. (71)–(73) plus (44) is deductive. Note that eq. (83) allows purely time-dependent healing to occur significantly when $\Omega\ll\beta$, even in VS tests. Hence, strength evolution following a large negative step with $V_{\rm after}/V_{\rm before}\ll\beta$ is not symmetric with that following a positive step. This deviation from the canon is a side effect of the aging-slip switch during VS tests discussed in §4.1, also seen in the composite law for $V_{\rm after}<V_{\rm c}$. However, there is no experimental data for extremely small $V_{\rm after}/V_{\rm before}$, which is above $10^{-4}$ so far, so there is no evidence that the asymmetric behavior predicted for very large velocity steps is unfactual.
We now look at the roles of $\beta$ in terms of reproducing canonical behaviors. A function $f$ satisfying eqs. (79) and (83) reduces to $f\simeq f_{\rm VS}$ for $\Omega\gg\beta$ (from eqs. 45 and 72), while being $f\simeq f_{\rm SHS|NSC}$ for $\Omega\ll\beta$ (from eqs. 59 and 63). In other words, eqs. (79) and (83) divide the $\Omega$ range into the two canonical regimes of VS and SHS||||NSC:
@@
Thus, $\beta$ is observable as the scale of $\Omega$ at which the state evolution is switched between the VS and SHS||||NSC canons. Because the canon-constrained $f_{\rm VS}$ (eq. 72) is the same as the slip law (Ruina, 1983) and the $f_{\rm SHS|NSC}$ is the same as the aging law at $\Omega\ll 1$ ($\supset\Omega\ll\beta\ll 1$), eq. (84) also reads as the switching between the aging and slip laws according to $\Omega$. Moreover, eq. (84) elucidates how the unavoidable side effects of the aging-slip switch (§4.1) appear in our specific approach using the functional form of $f=\psi(\Omega)$ as well as the TDF $f=f_{t}+f_{\delta}V$; sufficiently below the steady state, the side effects harm the reproduction of the VS canon, and in the other ranges of $\Omega$ they harm the reproduction of the SHS canon. As will be discussed at the end of §4.2 (and also discussed from another viewpoint in §5.1), these side effects are commonly seen in evolution laws trying to reproduce the three canons simultaneously. We may say that these side effects are nearly a logical consequence of those canonical behaviors.
@@
As mentioned later (§5.3), Bhattacharya et al. (2022) concluded $|f-f_{t}|>|f-f_{\delta}V|$ for $\Omega$ down to $10^{-3.5}$, based on their VS tests with $V_{\rm after}/V_{\rm before}\geq 10^{-3.5}$. In contrast, also as mentioned there, Bhattacharya et al. (2015) showed that a VS test by Bayart et al. (2006) with $V_{\rm after}/V_{\rm before}=10^{-2}$ indicates that the strength recovery following this large negative step is slightly faster than the slip-law prediction with parameters determined from VS tests with $V_{\rm after}/V_{\rm before}\geq 10^{2}$, $10^{1}$, $10^{-1}$, which suggests the transition from the slip law to the aging law starts kicking in at $\Omega=10^{-2}$. They concluded that this slight behavior difference for $V_{\rm after}/V_{\rm before}=10^{-2}$ is statistically significant through the significantly different $c$-value derived from the fitting by Nagata’s law, as detailed in §5.3. We leave the range $10^{-3.5}<\Omega<10^{-2}$ as a grey zone for now. It is natural to expect that $\beta$ depends on the details of the experimental conditions. In the rest of the paper, we assume $\beta\ll 1$.
@@ 4.2.2 Derivation without heuristics
So far, we have taken a strategy to search evolution functions under our heuristic ansatz that presumes unary functions of $\Omega$ (eq. 77, or equivalently  78) so that the SS canon is secured beforehand. Hence, in deriving eq. (83), the constraint on $f_{t}$ ($=\psi_{\rm c}(\Omega)$ by the ansatz), we only had to consider VS and SHS constraints. In §4.2.2, we will derive constraint on $f_{t}$ without any heuristics. The resultant constraint on $f_{t}$ is a sufficient and necessary condition to reproduce the three canons, whereas eq. (83), derived under the heuristic ansatz, was a sufficient condition. Note that the following argument holds within the standard rate and state framework of $\dot{\theta}=f(V,\theta)=f_{t}+f_{\delta}V$, with $f_{\delta}V$ having the form of eq. (72) directly constrained by the VS canon (plus a part of SS constraint, strictly speaking); then $f_{t}$ considered in the following argument is $f_{t}(V,\theta)$. These are the same circumstances as in the derivation of eq. (83), except that the unary ansatz has been removed.
We begin with the VS and SHS canons. In the introductory part of §4, we have shown that the NSC ($|f_{t}|\gg|f_{\delta}V|$) holds at $\Omega\to 0$, and the NSC can exist only within a right-bounded interval of $\Omega$ [$\Omega\ll D_{\rm c}/L$ for $|\ln D_{\rm c}/L|\lesssim 1$; for $|\ln D_{\rm c}/L|\gg 1$, $\Omega\ll D_{\rm c}/L/|\ln(D_{\rm c}/L)|$]. Thus, the existence of the NSC sets a lower bound of the $\Omega$ value sufficiently below which $|f_{t}|\gg|f_{\delta}V|$ must hold. On the other hand, the VS canon imposes the opposite, namely $|f_{t}|\ll|f_{\delta}V|$ (eq. 44) within a left-bounded interval of $\Omega$: $\Omega\geq V_{\rm after}/V_{\rm before}(\ll 1)$, which is experienced during the state evolution following a negative velocity step, for any $V=V_{\rm after}$. That is, $|f_{t}|\ll|f_{\delta}V|$ in VS tests sets an upper bound of the $\Omega$ value sufficiently above which $|f_{t}|\ll|f_{\delta}V|$. Hence, to satisfy these $\Omega$ conditions to realize $|f_{t}|\gg|f_{\delta}V|$ and $|f_{t}|\ll|f_{\delta}V|$, there must exist an $\Omega$ value $\beta$ such that $0<\beta\ll V_{\rm after}/V_{\rm before}(\ll 1)$, sufficiently below which $|f_{t}|\gg|f_{\delta}V|$ and sufficiently above which $|f_{t}|\ll|f_{\delta}V|$. Accordingly, $f_{t}$ is constrained as follows: $f_{t}$ is much smaller than $f_{\delta}V$ (eq. 72, from VS) for $\Omega\gg\beta$, while it must conform $f_{t|{\rm NSC}}$ (eq. 73, from SHS||||NSC) for $\Omega\ll\beta$:
@@
Equation (86) differs from eq. (83) only in that eq. (86) constrains $f_{t}$ rather than $\psi_{\rm c}$; the right-hand side, i.e., the required dependency on $V$ and $\theta$, is the same. This stands to reason because the derivation was based on SHS||||NSC and VS canons, as in the derivation of eq. (83). We may regard eq. (86) as the generalized version of eq. (83).
Now, we consider the SS canon, which seems no longer guaranteed beforehand because $f_{t}$ considered now is not limited to unary functions of $\Omega$. However, as shown below, the SS canon is achieved without additionally imposing $\Omega_{\rm SS}=1$ (eq. 71); $f_{t}$ satisfying eq. (86) vanishes near the steady state so that it does not affect the steady-state value $\Omega=1$ set by $f_{\delta}V$. From eqs. (72) and  (86), the evolution function becomes $f=-(L/D_{\rm c})\Omega\ln\Omega+o[(L/D_{\rm c})\Omega\ln\Omega]$ for $\Omega\gg\beta$, which already guarantees that $\Omega=1$ is a fixed point of the evolution law. Furthermore, if a small additional requirement of non-negativity of $f_{t}$ is imposed around $\Omega\sim\beta$ (note eq. (86) does not constrain $f_{t}$ for this $\Omega$ range),
@@
Thus, the above derived set of constraints eqs. (72), (86), and (87), where the heuristic ansatz has been removed, constitute a necessary and sufficient condition to satisfy the experimental constraints 1–3 (eqs. 7172) and auxiliary constraint (eq. 44), that is, to reproduce all the canonical behaviors (except for the $\Omega\sim\beta$, the border regime of the aging-slip switch).
@@
The above argument elucidates an interesting point about the role of eq. (83). In deriving eq. (83), we intended to secure the SS canon by the unary ansatz. However, it has turned out that any function having the asymptotic $V$, $\theta$ dependency conforming eq. (86), practically the same as eq. (83) if not unary functions of $\Omega$, automatically conforms the SS canon under the trivial assumption of non-negative $f_{t}$. That is, retrospectively speaking, the ansatz was not really necessary for our original aim of securing the SS canon. Equations (83) and (86) both state that $f_{t}$ must be negligible unless the system is sufficiently below the steady state, which is why the steady state is unchanged from the fixed point of $f_{\delta}V$, $\Omega_{\rm SS}=1$. Nevertheless, in the subsequent development (§4.34.5) we basically keep the unary ansatz and use eq. (83), rather than eq. (86). However, we now have to say that the substantial role of the ansatz is a mere pragmatic convention to avoid considering too many candidates in proposing evolution laws.
@@
In any case, with or without the unary ansatz, the aging-slip switch according to $\Omega$ (eq. 83, or eq. (86) thus follows from the three canons, as long as $\dot{\theta}=f(V,\theta)$. At the end of this subsection, we illustrate this logical consequence by compiling relevant VS and SHS||||NSC tests (Fig. 9). Given the constraints considered here (eqs. 31, 43, and 65, plus eq. 44), the possible functional form of $f$ is limited to one of the following two: $f\simeq f_{\delta}V=-(L/D_{\rm c})\Omega\ln\Omega$, which reproduces the VS canon, or $f\simeq f_{t|{\rm NSC}}(=1)$, which reproduces the SHS||||NSC canon. Hence, the remaining task is to identify, in the $(V,\theta)$ space, the regime where evolution agrees with the aging law but contradicts the slip law (blue and pale blue areas in Fig. 9) and the regime where evolution agrees with the slip law but contradicts the aging law (red and pale red areas in Fig. 9).
[…]
@@
Figure 9: Colored map of the $(V,\theta)$ regimes where the aging law (blue) is required and the slip law (red) is required based on the compilation of existing experiments. $V_{*}=1$ $\mu$m/s is here assumed. Pale-colored regimes represent where those behaviors are expected; though the experiments have not probed those regimes, we think of no particular reason to expect the behaviors to change. The red regime is based on Kilgore et al. (1993) and Bhattacharya et al. (2015), while the blue regime is based on Nakatani & Mochizuki (1996). See the text for detailed explanations.

### v3

@@ 4.1 Strategy to reconcile the aging law and the slip law
However, a well-known side effect exists in its prediction of steady-state $\Phi_{\rm SS}(V)$, as pointed out by Kato & Tullis (2001) themselves (Fig. 6). As shown in Fig. 6b, $\Phi_{\rm SS}(V)$ for $V\ll V_{\rm c}$ is predicted to be higher than the experimental requirement (eq. 26) by $W(1)B$, where $W(1)=0.567...$ is the omega constant, contradicting the SS canon. This level shift of $\Phi_{\rm SS}$ around $V_{\rm c}$ results from the extra strengthening mechanism (i.e., the aging law) operating at $V\ll V_{\rm c}$, in addition to the slip law, which alone matches eq. (26) exactly. By fitting $\tau_{\rm SS}(V)$ data of Blanpied et al. (1998b), Kato & Tullis (2001) suggested the possible level shift at $V_{\rm c}=0.01$ $\mu$m/s. However, the data have large scatterings and do not make strong evidence for the reality of the level shift. Furthermore, Kato & Tullis (2001) noted that $V_{\rm c}=0.01$ $\mu$m/s contradicts behaviors in VS tests, as explained in the next paragraph.
Figure 6: Comparison of steady-state values predicted by the composite law (blue) and those given by the SS constraint (eq. 71; red). Plotted against the slip rate $V$ normalized by $V_{\rm c}$. (a) Steady-state values of $\Omega=V\theta/L$. (b) Steady-state values of strength $\Phi$. Difference of $\Phi_{\rm SS}(V)$ from the canonical value $\Phi_{*}-B\ln(V/V_{*})$, given by eqs. (11) and (71) with the reference slip rate $V_{*}$ taken so as to satisfy $V_{*}\gg V_{\rm c}$, is plotted after being normalized by $B$. $\Delta\Phi_{\rm SS}=\Phi_{\rm SS}(V)-\Phi_{\rm SS}(V_{*})$ denotes the variation in $\Phi_{\rm SS}$.
@@
Lastly, we briefly describe our strategy to reproduce the three canons. Among the three defects of the composite law mentioned above, the level shift of $\Phi_{\rm SS}(V)$, which violates the SS canon (eq. 26), is not directly rooted in the antinomy between the VS constraint (eq. 72) and SHS||||NSC constraint (eq. 73). Hence, the present development aims to predict $\Phi_{\rm SS}(V)$ exactly as eq. (26). In §4.2, we will show that it is possible to construct evolution laws that perfectly match the SS canon (eq. 26) for all the ranges of $V$. Moreover, we will find functional forms to do so are fairly restricted. This finding has facilitated our development. Within such functional forms, we will try to narrow the conditions where the other two unavoidable deviations from canons during the strengthening phase of VS and SHS||||NSC tests arise. Eventually, we could see that our proposed evolution laws violate canons only at extreme conditions outside the conditions experimentally well verified so far. As the constraints eqs. (72) and (73) are the slip law and the aging law for $\Omega\ll 1$, respectively, the present development, in which we purposefully control the unavoidable side effects of the aging-slip switch, squarely answers the long-standing problem of how to reconcile these two laws (Ampuero & Rubin, 2008).
Figure 7: Strength evolution in ideal VS tests predicted by the composite law. $L=D_{\rm c}$ and $\theta\gg\theta_{\rm X}$ are assumed. $V_{\rm before}$ is fixed at $100V_{\rm c}$. The visualizing scheme is the same as in Fig. 2. Solid lines of different colors show the strength evolutions for different values of $V_{\rm after}$. The curves of $V_{\rm after}/V_{\rm before}=10^{-3}$ and $10^{-4}$ nearly overlap with each other due to the aging-law-like blowups of the early stage. As a reference, the canonical behavior (eq. 32) for each $V_{\rm after}$ is shown by the dotted line of the same color, although not visible in cases with $V_{\rm after}\geq 0.1V_{\rm before}$ because it almost coincides with the prediction by the composite law.
@@ 4.2.1 Heuristic derivation
As $\psi$ is a rewritten form of $\phi$, $\psi$ has a unique zero at $\Omega=C_{\rm SS}$ [$\psi(\Omega)=0\Leftrightarrow\Omega=C_{\rm SS}$]. Figure 8 shows the behavior of evolution laws given by eq. (77) (or equivalently, eq. 78). In the $V$-$\theta$ space, $\dot{\theta}$ varies only along the direction of the $V\theta$-axis while remaining constant along the direction normal to it (dotted lines in Fig. 8a).
@@
For $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, the expressions of $f_{t|{\rm NSC}}$ and $f_{\delta}V$ (eqs. 72 and 73) both fit the functional form eq. (78) we consider. As far as eq. (78) holds, the ratio $f_{t}/f_{t|{\rm NSC}}$ is also a function of $\Omega$, which we denote by $\psi_{\rm c}(\Omega)$ hereafter. Then, under eqs. (72), (73), and (78), the evolution function $f$ under the TDF (eq. 25) can be expressed as
@@
So far, eq. (77) is the only heuristics. Otherwise, the derivation down to eq. (83) from eqs. (71)–(73) plus (44) is deductive. Note that eq. (83) allows purely time-dependent healing to occur significantly when $\Omega\ll\beta$, even in VS tests. Hence, strength evolution following a large negative step with $V_{\rm after}/V_{\rm before}\ll\beta$ is not symmetric with that following a positive step. This deviation from the canon is a side effect of the aging-slip switch during VS tests discussed in §4.1, also seen in the composite law for $V_{\rm after}<V_{\rm c}$. However, currently there are no experimental data for $V_{\rm after}/V_{\rm before}$ less than $10^{-3.5}$ (Bhattacharya et al., 2022), so there is no evidence that the asymmetric behavior predicted for such extreme velocity steps is unfactual.
We now look at the roles of $\beta$ in terms of reproducing canonical behaviors. A function $f$ satisfying eq. (79) reduces to $f\simeq\psi_{\rm c}f_{\rm SHS|NSC}+f_{\rm VS}$ (given eqs. 45 and 46), leading to $f\simeq f_{\rm VS}$ for $\Omega\gg\beta$ and $f\simeq f_{\rm SHS|NSC}$ for $\Omega\ll\beta$ from eq. (83). In other words, eqs. (79) and (83) divide the $\Omega$ range into the two canonical regimes of VS and SHS||||NSC:
@@
Thus, $\beta$ is observable as the scale of $\Omega$ at which the state evolution is switched between the VS and SHS||||NSC canons. Because the canon-constrained $f_{\rm VS}$ (eq. 72) is the same as the slip law (Ruina, 1983) and the $f_{\rm SHS|NSC}$ is the same as the aging law for $\Omega\ll 1$ ($\supset\Omega\ll\beta$), eq. (84) also reads as the switching between the aging and slip laws according to $\Omega$. Moreover, eq. (84) elucidates how the unavoidable side effects of the aging-slip switch (§4.1) appear in our specific approach using the functional form of $f=\psi(\Omega)$ as well as the TDF $f=f_{t}+f_{\delta}V$; sufficiently below the steady state, the side effects harm the reproduction of the VS canon, and in the other ranges of $\Omega$ they harm the reproduction of the SHS canon. As will be discussed at the end of §4.2 (and also discussed from another viewpoint in §5.1), these side effects are commonly seen in evolution laws trying to reproduce the three canons simultaneously. We may say that these side effects are nearly a logical consequence of those canonical behaviors.
@@
As mentioned later (§5.3), Bhattacharya et al. (2022) concluded $|f-f_{t}|>|f-f_{\delta}V|$ for $\Omega$ down to $10^{-3.5}$, based on their VS tests with $V_{\rm after}/V_{\rm before}\geq 10^{-3.5}$. In contrast, also as mentioned there, Bhattacharya et al. (2015) showed that a VS test by Bayart et al. (2006) with $V_{\rm after}/V_{\rm before}=10^{-2}$ indicates that the strength recovery following this large negative step is slightly faster than the slip-law prediction with parameters determined from VS tests with $V_{\rm after}/V_{\rm before}\geq 10^{2}$, $10^{1}$, $10^{-1}$, which suggests the transition from the slip law to the aging law starts kicking in at $\Omega=10^{-2}$. They concluded that this slight behavior difference for $V_{\rm after}/V_{\rm before}=10^{-2}$ is statistically significant, though the slight difference may not be sufficient to affirm an extra parameter to express the Aging-Slip switch, as detailed in §5.3. We leave the range $10^{-3.5}<\Omega<10^{-2}$ as a grey zone for now. It is natural to expect that $\beta$ depends on the details of the experimental conditions. In the rest of the paper, we assume $\beta\ll 1$.
@@ 4.2.2 Derivation without heuristics
So far, we have taken a strategy to search evolution functions under our heuristic ansatz that presumes unary functions of $\Omega$ (eq. 77, or equivalently,  78) so that the SS canon is secured beforehand. Hence, in deriving eq. (83), the constraint on $f_{t}$ ($=\psi_{\rm c}(\Omega)$ by the ansatz), we only had to consider VS and SHS constraints. In this §4.2.2, we derive constraint on $f_{t}$ without heuristics. The resultant constraint on $f_{t}$ is a sufficient and necessary condition to reproduce the three canons, whereas eq. (83), derived under the heuristic ansatz, was a sufficient condition. Note that the following argument holds within the standard rate-and-state framework of $\dot{\theta}=f(V,\theta)=f_{t}+f_{\delta}V$, with $f_{\delta}V$ having the form of eq. (72) directly constrained by the VS canon (plus part of the SS constraint, strictly speaking); then $f_{t}$ considered in the following argument is $f_{t}(V,\theta)$. These are the same circumstances as in the derivation of eq. (83), except that the unary ansatz has been removed.
We begin with the VS and SHS canons. In the introductory part of §4, we have shown that the NSC ($|f_{t}|\gg|f_{\delta}V|$) holds at $\Omega\to 0$, and the NSC can exist only within a right-bounded interval of $\Omega$ [$\Omega\ll D_{\rm c}/L$ for $|\ln D_{\rm c}/L|\lesssim 1$; for $|\ln D_{\rm c}/L|\gg 1$, $\Omega\ll D_{\rm c}/L/|\ln(D_{\rm c}/L)|$]. Thus, the existence of the NSC sets a lower bound of the $\Omega$ value sufficiently below which $|f_{t}|\gg|f_{\delta}V|$ must hold. On the other hand, the VS canon imposes the opposite, namely $|f_{t}|\ll|f_{\delta}V|$ (eq. 44) within a left-bounded interval of $\Omega$: $\Omega\geq V_{\rm after}/V_{\rm before}(\ll 1)$, which is experienced during the state evolution following a negative velocity step, for any $V=V_{\rm after}$. That is, $|f_{t}|\ll|f_{\delta}V|$ in VS tests sets an upper bound of the $\Omega$ value sufficiently above which $|f_{t}|\ll|f_{\delta}V|$. For these $\Omega$ ranges where $|f_{t}|\gg|f_{\delta}V|$ and $|f_{t}|\ll|f_{\delta}V|$, respectively, to coexist, there must exist an $\Omega$ value $\beta$ such that $0<\beta\ll V_{\rm after}/V_{\rm before}(\ll 1)$, sufficiently below which $|f_{t}|\gg|f_{\delta}V|$ and sufficiently above which $|f_{t}|\ll|f_{\delta}V|$. Accordingly, $f_{t}$ is constrained as follows: $f_{t}$ is much smaller than $f_{\delta}V$ (eq. 72, from VS) for $\Omega\gg\beta$, while it must conform $f_{t|{\rm NSC}}$ (eq. 73, from SHS||||NSC) for $\Omega\ll\beta$:
@@
Equation (86) differs from eq. (83) only in that eq. (86) constrains $f_{t}$ rather than $\psi_{\rm c}$; the right-hand side, i.e., the required dependency on $V$ and $\theta$, is the same. This stands to reason because the derivation was based on the SHS||||NSC and VS canons, as in the derivation of eq. (83). We may regard eq. (86) as the generalized version of eq. (83).
Now, we consider the SS canon, which seems no longer guaranteed beforehand because $f_{t}$ considered now is not limited to unary functions of $\Omega$. However, as shown below, the SS canon is achieved without additionally imposing $\Omega_{\rm SS}=1$ (eq. 71); $f_{t}$ satisfying eq. (86) vanishes near the steady state so that it does not affect the steady-state value $\Omega=1$ set by $f_{\delta}V$. From eqs. (72) and  (86), the evolution function becomes $f=-(L/D_{\rm c})\Omega\ln\Omega+o[(L/D_{\rm c})\Omega\ln\Omega]$ for $\Omega\gg\beta$, which already guarantees that $\Omega=1$ is a fixed point of the evolution law. Furthermore, if a small additional constraint of non-negativity of $f_{t}$ is imposed around $\Omega\sim\beta$ (note eq. 86 does not constrain $f_{t}$ for this $\Omega$ range):
@@
Thus, the above-derived set of constraints eqs. (72), (86), and (87), where the heuristic ansatz has been removed, constitute a necessary and sufficient condition to satisfy the experimental constraints 1–3 (eqs. 7173) and the auxiliary constraint (eq. 44), that is, to reproduce all the canonical behaviors (except for $\Omega\sim\beta$, the border regime of the aging-slip switch).
@@
The above argument elucidates an interesting point about the role of eq. (83). In deriving eq. (83), we intended to secure the SS canon by the unary ansatz. However, it has turned out that any function having the asymptotic $V$, $\theta$ dependency conforming eq. (86), practically the same as eq. (83), automatically conforms to the SS canon under the trivial assumption of non-negative $f_{t}$. That is, retrospectively speaking, the ansatz was in fact not necessary for our original aim of securing the SS canon. Both eqs. (83) and (86) state that $f_{t}$ must be negligible unless the system is sufficiently below the steady state, which is why the steady state is unchanged from the fixed point of $f_{\delta}V$, $\Omega_{\rm SS}=1$. Nevertheless, in the subsequent development (§4.34.5), we basically keep the unary ansatz and use eq. (83), rather than eq. (86). However, we now have to say that the substantial role of the ansatz is a mere pragmatic convention to avoid considering too many candidates in proposing evolution laws.
@@
In any case, with or without the unary ansatz, the aging-slip switch according to $\Omega$ (eq. 83 or 86) thus follows from the three canons, as long as $\dot{\theta}=f(V,\theta)$. At the end of this subsection, we illustrate this logical consequence by compiling relevant VS and SHS||||NSC tests (Fig. 9). Given the constraints considered here (eqs. 31, 43, and 65, plus eq. 44), the possible functional form of $f$ is limited to one of the following two: $f\simeq f_{\delta}V=-(L/D_{\rm c})\Omega\ln\Omega$, which reproduces the VS canon, or $f\simeq f_{t|{\rm NSC}}(=1)$, which reproduces the SHS||||NSC canon. Hence, the remaining task is to identify, in the $(V,\theta)$ space, the regime where evolution agrees with the aging law but contradicts the slip law (blue and pale-blue areas in Fig. 9) and the regime where evolution agrees with the slip law but contradicts the aging law (red and pale-red areas in Fig. 9).
Specifically, we used experiments of Kilgore et al. (1993) to define the velocity range ($10^{-4}$$10^{2}$ $\mu$m/s) of the red area (i.e., slip law required), where both VS and SS canons are required. The range of $\Omega$ where the VS canon is known required is the maximum of $|\log_{10}(V_{\rm after}/V_{\rm before})|$ for which the VS canon has been confirmed. Numerous experiments have confirmed the VS canon for up to 10-fold VS. Recently, Bhattacharya et al. (2015) have shown that the VS canon is observed even for 100-fold VS. Hence, we have decided that the VS canon has been confirmed up to 100-fold VSs and have adopted $10^{\pm 2}$ as the along-$\Omega$ axis ends of the red and pale-red bands. In addition, we have adopted $V_{\rm X}$ of about 100 $\mu$m/s from the high-speed VS tests at $\sigma=5$ MPa by Kilgore et al. (1993) because we will use SHS||||NSC tests at $\sigma=5$ MPa (Nakatani & Mochizuki, 1996) to define the $(V,\theta)$ range where the agreement with the aging-law-like (i.e., the canonical SHS||||NSC) behavior and the contradiction to the slip-law-like behavior have been confirmed. The pale-red area below $V=10^{-4}$ $\mu$m/s, the minimum slip rate of Kato & Tullis (2001), is included by extrapolation given we have no particular observation suggesting the VS behavior to change below a certain slip rate.
@@
Figure 9: Colored map showing the $(V,\theta)$ regimes where the aging law (blue) is required and the slip law (red) is required based on the compilation of existing experiments. $V_{*}=1$ $\mu$m/s is assumed. Pale-colored regimes represent where those behaviors are expected; though the experiments have not probed those regimes, we think of no particular reason to expect the behaviors to change. The red regime is based on Kilgore et al. (1993) and Bhattacharya et al. (2015), while the blue regime is based on Nakatani & Mochizuki (1996). See the text for details.